Potassium hydroxide reacts completely with sulfuric acid according to the equation:

H2SO4(aq) + 2KOH(aq) K2SO4(aq) + 2H2O(l)

0.95 g of potassium hydroxide were dissolved in water and made up to 250 cm3 of
solution. 25 cm3 samples of this solution were titrated against sulfuric acid (H2SO4)solution. The mean volume (titre) of sulfuric acid required for complete reaction was15.00 cm3.

(i) Find the molar concentration of the prepared solution of potassium hydroxide.

Complete the calculation given below:

Relative Molecular Mass (RMM) of

KOH

=

= ………………………… (1 mark)

No. of moles of KOH in 250 cm3 =

= ………………………… (1 mark)

 molar concentration of KOH =

(Include units in your answer) = ………………………… (1 mark)

Continued…..

19.00 cm3

20.00 cm3

21.00 cm3

Answer: …………………………………cm3
1
Expert's answer
2021-07-17T04:44:57-0400

Relative Molecular Mass (RMM) of

KOH is 56.11 g/mole


No. of moles of KOH in 250 cm3"= \frac{0.95}{56.11}=0.0169moles"


Molarity = moles/volume


Molarity = "\frac{0.0169}{56.11}=0.0003M"


Average ="\frac{19.00+20.00+21.00}{3}=20.00 cm^3"